OpenStudy (anonymous):

Find the general solution of each differential equation: dy/dx = (2x)/(e^2y)

OpenStudy (ribhu):

(e^2y)dy = 2x dx e^2y / 2 = x^2 +C

OpenStudy (ribhu):

second step obtained when 1st was integrated.

OpenStudy (anonymous):

Right I've gotten that far. What's next? Or is that where you end it?

OpenStudy (jhannybean):

Considering its separable, and you have integrated it, you're done

OpenStudy (anonymous):

Thanks! Sorry for posting the same question twice Ribhu!

OpenStudy (ribhu):

yeah this is teh way

OpenStudy (ribhu):

its fine u should understand it